Grade 12th passMechanics

(1)a particle moves in a straight line with constant speed 4m/s for 2sec. then with 6m/s for 3sec. find the av.speed of the particle in the given time interval.

sanjeev rana
8 Years agoGrade 12th pass

2 Answers

sreehari
8 Years ago
Distance Travelled = Speed x Time taken.
Avg Speed = (Total Distance travelled)/(Total Time taken)
In (1);
    Distance Travelled = 4 x 2 = 8 m.
In (2);
    Distance Travelled = 6 x 3 = 18 m.
 
Total Distance Travelled = 26 m.
Hence, Average speed = 26/5 = 5.2 m/sec
Adarsh
5 Years ago
Speed= distance/time so in the first case 
4=d1/2 so d1= 8m
In the second case using speed formula
6=d2/3 so d2=18m
Total distance= d1+d2=18+8=26
Total time =2+3=5.        (Time is given in the question)
Average speed= total distance/total time= 26/5 =5.2 m/s